# Correcting the Mythos of KL-Regularization: Direct Alignment without Overoptimization via $\chi^{2}$-Preference Optimization

### v2

@@ Abstract
Language model alignment methods, such as reinforcement learning from human feedback (RLHF), have led to impressive advances in language model capabilities, but existing techniques are limited by a widely observed phenomenon known as overoptimization, where the quality of the language model plateaus or degrades over the course of the alignment process. Overoptimization is often attributed to overfitting to an inaccurate reward model, and while it can be mitigated through online data collection, this is infeasible in many settings. This raises a fundamental question: Do existing offline alignment algorithms make the most of the data they have, or can their sample-efficiency be improved further?
We address this question with a new algorithm for offline alignment, $\chi^{2}$-Preference Optimization ($\chi$PO).
$\chi$PO is a one-line change to Direct Preference Optimization (DPO; Rafailov et al. (2023)), which only involves modifying the logarithmic link function in the DPO objective. Despite this minimal change, $\chi$PO implicitly implements the principle of pessimism in the face of uncertainty via regularization with the $\chi^{2}$-divergence—which quantifies uncertainty more effectively than KL-regularization—and provably alleviates overoptimization, achieving sample-complexity guarantees based on single-policy concentrability—the gold standard in offline reinforcement learning. $\chi$PO’s simplicity and strong guarantees make it the first practical and general-purpose offline alignment algorithm that is provably robust to overoptimization.
@@ 1 Introduction
Alignment methods like RLHF have led to significant advances in language model capabilities, particularly in chat domains, but existing techniques are limited by a widely observed phenomenon known as reward overoptimization or reward hacking (Michaud et al., 2020; Tien et al., 2022; Gao et al., 2023; Rafailov et al., 2024a). Since the reward model used implicitly or explicitly by these methods is an imperfect proxy for human preferences, the true quality of the language model can degrade as training proceeds, even as performance under the reward model continues to improve. Intuitively, this occurs because the language model may drift away from the manifold covered by the human-labeled data used to train the reward model and end up in a region where the reward model is inaccurate.
@@
  • Is overoptimization in offline alignment an information-theoretic phenomenon? This would mean that there is simply not enough information in the human-labeled (offline) preference dataset due to partial coverage, and no algorithmic intervention (e.g., early stopping), no matter how clever, can avoid the overoptimization issue.
@@
Parallel developments in the theory of offline reinforcement learning suggest that the answer may be the latter. Indeed, this literature has addressed the challenge of overoptimization—typically referred to as distribution shift—through the principle of pessimism in the face of uncertainty. The principle of pessimism asserts that given an offline dataset with partial coverage, a decision maker should choose their response according to the most pessimistic view of the world supported by the data. Pessimism encourages the model to avoid overfitting to the offline dataset and is supported by a rich theory, showing that it offers provable robustness to overoptimization in stylized settings (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021).
Perhaps the greatest barrier to implementing pessimism for language models is to efficiently quantify uncertainty in the learned reward model and distill this understanding into an actionable form. This is particularly challenging due to the combinatorial nature of the space of responses. A natural approach that forms the basis of most existing offline alignment methods (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022; Rafailov et al., 2023) is to use KL-regularization to encourage the learned policy to stay close to a reference policy (typically obtained through supervised fine-tuning and used to collect the offline data), but this form of regularization is insufficient to induce pessimism (Gao et al., 2023), and is provably suboptimal in theory (Zhu et al., 2023; Song et al., 2024, see also Section A.1). On the other hand, offline reinforcement learning theory offers abstract pessimistic algorithms that are suitable—at least statistically—for large models (Xie et al., 2021; Uehara and Sun, 2021; Zhan et al., 2022; Chen and Jiang, 2022), but cannot be implemented directly without losing fidelity to the theory or making unrealistic modeling assumptions (Zhu et al., 2023; Zhan et al., 2023b; Li et al., 2023; Xiong et al., 2023; Liu et al., 2024; Cen et al., 2024; Fisch et al., 2024; Ji et al., 2024). Notably, we show (Section A.1) that the so-called “DPO+SFT” approach developed by Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) requires a certain unrealistic convexity property and is provably suboptimal in its absence. Can we develop practical offline alignment methods with provable robustness to overoptimization by exploiting the unique structure of the language modeling problem?
@@ 1.1 Contributions
Our analysis of $\chi$PO leverages several new techniques. First, RLHF with $\chi^{2}$-regularization is sufficient to achieve guarantees based on single-policy concentrability (Section 3.1 and Appendix C). Next, we show that a variant of the DPO reparameterization trick that combines $\chi^{2}$-regularization with KL-regularization (“mixed” $\chi^{2}$-regularization) can be used to reformulate our objective into a purely policy-based objective, in spite of the fact that $\chi^{2}$-regularization fails to satisfy certain regularity conditions found in prior work (Wang et al., 2023a). Finally, and perhaps most importantly, we use a novel analysis to show that pessimism is preserved after reparameterization.
@@ 1.2 Paper Organization
Section 2 provides background on offline alignment and the suboptimality of existing algorithms. Section 3 presents our main algorithm, $\chi$PO, and accompanying theoretical guarantees. Section 4 then presents detailed intuition into how $\chi$PO modulates the bias-overoptimization tradeoff and implements pessimism, and Section 5 sketches the proof for its main statistical guarantee.
Section 6 contains results for general preference models, including an impossibility result for obtaining guarantees under single-policy concentrability in this setting. We conclude with discussion in Section 7. Proofs and additional results are deferred to the appendix, with highlights including (i) detailed discussion on suboptimality of existing pessimistic approaches (Appendix A), and (ii) additional algorithms and guarantees based on the $\chi^{2}$-regularization framework (Appendix C).
@@ 2.1 Alignment from Human Feedback
Following prior work (e.g., Rafailov et al. (2023); Ye et al. (2024)), we adopt a contextual bandit formulation of the alignment problem. We formalize the language model as a policy $\pi:\mathcal{X}\to\Delta(\mathcal{A})$ which maps a context (prompt) $x\in\mathcal{A}$ to an action (response) $a\in\mathcal{A}$ via $a\sim{}\pi(\cdot\mid{}x)$, and let $\rho\in\Delta(\mathcal{X})$ denote the distribution over contexts/prompts.
@@
In the offline alignment problem (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022), we assume access to a dataset $\mathcal{D}_{\mathsf{pref}}=\left\{(x,a_{+},a_{-})\right\}$ of prompts and labeled response pairs generated from a reference policy (language model) $\pi_{\mathsf{ref}}$; the reference policy is typically obtained through supervised fine tuning. Here, $a_{+}$ is a positive action/response and $a_{-}$ is a negative action/response. Given the context/prompt $x\sim\rho$, the pair $(a_{+},a_{-})$ is generated by sampling a pair $(a,b)$ as $a\sim{}\pi_{\mathsf{ref}}(\cdot\mid{}x)$ and $b\sim{}\pi_{\mathsf{ref}}(\cdot\mid{}x)$, and then ordering them as $(a_{+},a_{-})$ based on a binary preference $y\sim{}\mathbb{P}(a\succ{}b\mid{}b)$. We assume that preferences follow the Bradley-Terry model (Bradley and Terry, 1952), in which
@@
for an unknown reward function $r^{\star}:\mathcal{X}\times\mathcal{A}\to\left[0,R_{\mathsf{max}}\right]$ for some $R_{\mathsf{max}}\geq 1$. Based on the preference dataset $\mathcal{D}_{\mathsf{pref}}$, we aim to learn a policy $\widehat{\pi}$ that has high reward in the sense that
$\max_{\pi}J(\pi)-J(\widehat{\pi})\leq\varepsilon,$
for a small $\varepsilon>0$, where $J(\pi)\vcentcolon={}\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x)}\left[r^{\star}(x,a)\right]$. Going forward, we abbreviate $\operatorname{\mathbb{E}}_{\pi}\left[\cdot\right]\vcentcolon={}\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x)}[\cdot]$. Without loss of generality, we assume that $\rho(x)>0$ for all $x$ and $\pi_{\mathsf{ref}}(a\mid x)>0$ for all $x,a$.
@@
where we adopt the shorthand $D_{\mathsf{KL}}\left(\pi\,\|\,\pi^{\prime}\right)=\operatorname{\mathbb{E}}_{x\sim\rho}\left[D_{\mathsf{KL}}\left(\pi(\cdot\mid{}x)\,\|\,\pi^{\prime}(\cdot\mid{}x)\right)\right]$. These methods first estimate a reward function $\widehat{r}$ from $\mathcal{D}_{\mathsf{pref}}$ using maximum likelihood under the Bradley-Terry model, i.e.
@@
where $\sigma(x)\vcentcolon={}\frac{\exp(x)}{1+\exp(x)}$ is the sigmoid function and $\mathcal{R}$ is a class of reward functions, which is typically parameterized by a neural network. Then, they apply standard policy optimization methods like PPO to optimize the following estimated version of the KL-regularized objective:
@@ 2.2 Overoptimization and Insufficiency of KL-Regularization
Empirically, both classical RLHF and direct alignment methods like DPO have been observed to suffer from overoptimization (Gao et al., 2023; Guo et al., 2024; Rafailov et al., 2024a; Song et al., 2024), wherein model quality degrades during the optimization process as the learned policy drifts away from $\pi_{\mathsf{ref}}$. This can be mitigated by online alignment techniques (Gao et al., 2024; Guo et al., 2024; Dong et al., 2024; Xie et al., 2024), which collect labeled preference data on-policy during training, but there are many settings where this is impractical or infeasible. As we will see, the overoptimization phenomena in offline alignment methods is an issue of sample-inefficiency, which can be understood through the lens of coverage coefficients developed in the theory of offline reinforcement learning (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021).
@@
In the theory of offline reinforcement learning, the sample efficiency of offline algorithms is typically quantified through coverage coefficients (or, concentrability coefficients), which measure the quality of the data collected by the policy $\pi_{\mathsf{ref}}$ (Farahmand et al., 2010; Xie and Jiang, 2020; Zanette et al., 2021). Define the $L_{\infty}$-concentrability coefficient for a policy $\pi$ via
$\mathcal{C}_{\infty}^{\pi}\vcentcolon={}\sup_{(x,a)\in\mathcal{X}\times\mathcal{A}}\left\{\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right\}.$
For general offline RL, algorithms based on the principle of pessimism can achieve sample complexity guarantees that scale with the single policy concentrability coefficient $\mathcal{C}_{\infty}^{\pi^{\star}}$ for the optimal policy $\pi^{\star}$ (or more generally, any comparator policy), while non-pessimistic algorithms typically achieve guarantees based on the less favorable all-policy concentrability coefficient $\max_{\pi\in\Pi}\mathcal{C}_{\infty}^{\pi}$ (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021). Sample complexity guarantees scaling with single-policy concentrability reflect robustness to overoptimization, as they ensure that the algorithm has non-trivial sample complexity even if the data collection policy $\pi_{\mathsf{ref}}$ has poor coverage. In particular, the algorithm’s performance is never much worse than that of $\pi_{\mathsf{ref}}$ itself. On the other hand, an algorithm with performance depending on all-policy concentrability is susceptible to the overoptimization phenomena and can perform poorly when $\pi_{\mathsf{ref}}$ has poor coverage.
@@
For the stylized special case of alignment with linearly parameterized policies where $\pi_{\theta}(a\mid{}x)\propto{}\exp(\left\langle\phi(x,a),\theta\right\rangle)$ for a known feature embedding $\phi(x,a)\in\mathbb{R}^{d}$, Zhu et al. (2023) (see also Zhu et al. (2024); Song et al. (2024)) present analogous findings, highlighting that PPO and DPO are suboptimal with respect to dependence on the concentrability coefficient. They show that there exists an algorithm (based on pessimism) which for all policies $\pi^{\star}$ simultaneously achieves
@@
Meanwhile, PPO and DPO must suffer sample complexity scaling with $\max_{\pi}\mathcal{C}_{\infty}^{\pi}$. [^1] Zhu et al. (2023) achieve guarantees based on a feature coverage coefficient, which improves on concentrability coefficients, but is specialized to linear function approximation.
While the results above are encouraging, they are restricted to linearly parameterized policies, and cannot be directly applied to large language models. Most existing theoretical algorithms for offline alignment are similar in nature, and either place restrictive assumptions on the policy class $\Pi$ (Zhu et al., 2023; Zhan et al., 2023b; Li et al., 2023; Xiong et al., 2023) or are not practical to implement in a way that is faithful to theory (Ye et al., 2024; Ji et al., 2024).
@@
which augments the DPO objective (the second term) with an additional supervised fine-tuning-like (SFT) loss (the first term). While this objective is simple to apply to general policy classes, the existing single-policy concentrability guarantees for this method assume that $\Pi$ satisfies restrictive convexity conditions which do not hold in practice for large language models. Perhaps surprisingly, we show (Section A.1) that without convexity, the objective in Eq. 5 fails to achieve a single-policy concentrability guarantee. [^2] This finding is rather surprising because Xie et al. (2024) show that an optimistic counterpart to Eq. 5, which negates the SFT term, enjoys strong online RLHF guarantees with general policy classes without requiring analogous convexity conditions. In other words, DPO+SFT is insufficient to mitigate overoptimization.

### v3

@@ Abstract
Language model alignment methods such as reinforcement learning from human feedback (RLHF) have led to impressive advances in language model capabilities, but are limited by a widely observed phenomenon known as overoptimization, where the quality of the language model degrades over the course of the alignment process. As the model optimizes performance with respect to an offline reward model, it overfits to inaccuracies and drifts away from preferred responses covered by the data. To discourage such distribution shift, KL-regularization is widely employed in existing offline alignment methods, but overoptimization continues to harm performance. Lending theoretical insight into the source of these empirical observations, we first show that the KL-regularization is too weak to prevent overfitting, then raise the following question: is it possible to design an efficient algorithm that is provably robust to overoptimization?
We address this question with a new algorithm for offline alignment, $\chi^{2}$-Preference Optimization ($\chi$PO). $\chi$PO is a one-line change to Direct Preference Optimization (DPO; Rafailov et al. (2023)), which only involves modifying the logarithmic link function in the DPO objective. Despite this minimal change, $\chi$PO implicitly implements the principle of pessimism in the face of uncertainty via regularization with the $\chi^{2}$-divergence—which quantifies uncertainty more effectively than KL-regularization—and provably alleviates overoptimization, achieving sample-complexity guarantees based on single-policy concentrability—the gold standard in offline reinforcement learning. $\chi$PO’s simplicity and strong guarantees make it the first practical and general-purpose offline alignment algorithm that is provably robust to overoptimization.
@@ 1 Introduction
Alignment methods like RLHF have led to significant advances in language model capabilities, but existing techniques are limited by a widely observed phenomenon known as reward overoptimization or reward hacking (Michaud et al., 2020; Tien et al., 2022; Gao et al., 2023; Rafailov et al., 2024a). Since the reward model is an imperfect proxy for human preferences, the true quality of the language model can degrade as training proceeds, even as its performance under the reward model continues to improve. Intuitively, this occurs because the language model may drift away from the manifold covered by the human-labeled data used to train the reward model and end up in a region where the reward model is inaccurate.
@@
  • Is overoptimization in offline alignment an information-theoretic phenomenon? This would mean that there is simply not enough information in the human-labeled (offline) preference dataset due to partial coverage, and no algorithmic intervention can avoid the overoptimization issue.
@@
Previous developments in the theory of offline reinforcement learning suggest that the answer may be the latter. Indeed, this literature has addressed the challenge of overoptimization—typically referred to as distribution shift—through the principle of pessimism in the face of uncertainty, which asserts that, given an offline dataset with partial coverage, a decision maker should choose their response according to the most pessimistic view of the world supported by the data. Pessimism encourages the model to avoid overfitting to the offline dataset and is supported by a rich theory showing that it offersprovable robustness to overoptimization in stylized settings (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021).
Perhaps the greatest barrier to implementing pessimism in language models is the efficient quantification of uncertainty in the offline reward, and the distillation of this information into actionable form. Most existing offline alignment methods employ KL-regularization, which penalizes the learned policy for drifting from the reference policy, but this form of uncertainty quantification is insufficient to induce pessimism (Gao et al., 2023) and is provably suboptimal in theory (Zhu et al., 2023; Song et al., 2024, see also Section A.1). On the other hand, offline reinforcement learning theory offers abstract pessimistic algorithms that are suitable—at least statistically—for large models (Xie et al., 2021; Uehara and Sun, 2021; Zhan et al., 2022; Chen and Jiang, 2022), but cannot be implemented directly without losing theoretical fidelity or making unrealistic modeling assumptions (Zhu et al., 2023; Zhan et al., 2023a; Li et al., 2023; Xiong et al., 2023; Liu et al., 2024; Cen et al., 2024; Fisch et al., 2024; Ji et al., 2024). Notably, the so-called “DPO+SFT” approach developed by Liu et al. (2024); Cen et al. (2024); Fisch et al. (2024) is provably suboptimal unless the language model satisfies an unrealistic convexity property (Section A.1). Can we develop practical offline alignment methods with provable robustness to overoptimization by exploiting the unique structure of the language modeling problem?
@@ 1.1 Contributions
Our analysis of $\chi$PO leverages several new techniques. First, we show that RLHF with $\chi^{2}$-regularization is sufficient to achieve guarantees based on single-policy concentrability (Section 3.1 and Appendix B). Next, we show that a variant of the DPO reparameterization trick that combines $\chi^{2}$-regularization with KL-regularization (“mixed” $\chi^{2}$-regularization) can be used to reformulate our objective into a purely policy-based objective, in spite of the fact that $\chi^{2}$-regularization fails to satisfy certain regularity conditions found in prior work (Wang et al., 2023a). Finally, and perhaps most importantly, we use a novel analysis to show that pessimism is preserved after reparameterization.
@@ 1.2 Paper Organization
Section 2 provides background on offline alignment and the suboptimality of existing algorithms. Section 3 presents our main algorithm, $\chi$PO, and accompanying theoretical guarantees. Section 4 then presents detailed intuition into how $\chi$PO modulates the bias-overoptimization tradeoff and implements pessimism, and Section 5 sketches the proof for its main statistical guarantee. We perform experimental evaluations of $\chi$PO against DPO in the TL;DR summarization task (Stiennon et al., 2020), which is included in Section 6.
Section 7 contains results for general preference models, including an impossibility result for obtaining guarantees under single-policy concentrability in this setting. We conclude with discussion in Section 8. Proofs and additional results are deferred to the appendix, with highlights including (i) detailed discussion on suboptimality of existing pessimistic approaches (Appendix A), and (ii) additional algorithms and guarantees based on the $\chi^{2}$-regularization framework (Appendix B).
@@ 2.1 Alignment from Human Feedback
Following prior work (e.g., Rafailov et al. (2023); Ye et al. (2024)), we adopt a contextual bandit formulation of the alignment problem. We formalize the language model as a policy $\pi:\mathcal{X}\to\Delta(\mathcal{A})$ which maps a context (prompt) $x\in\mathcal{X}$ to an action (response) $a\in\mathcal{A}$ via $a\sim{}\pi(\cdot\mid{}x)$, and let $\rho\in\Delta(\mathcal{X})$ denote the distribution over contexts/prompts.
@@
In the offline alignment problem (Christiano et al., 2017; Bai et al., 2022; Ouyang et al., 2022), we assume access to a dataset $\mathcal{D}_{\mathsf{pref}}=\left\{(x,a_{+},a_{-})\right\}$ of $n$ prompts and labeled response pairs generated from a reference policy (language model) $\pi_{\mathsf{ref}}$, which is typically obtained through supervised fine tuning. Here, $a_{+}$ is a positive action/response and $a_{-}$ is a negative action/response. Given the context/prompt $x\sim\rho$, the pair $(a_{+},a_{-})$ is generated by sampling a pair $(a,b)$ as $a\sim{}\pi_{\mathsf{ref}}(\cdot\mid{}x)$ and $b\sim{}\pi_{\mathsf{ref}}(\cdot\mid{}x)$, and then ordering them as $(a_{+},a_{-})$ based on a binary preference $y\sim{}\mathbb{P}(a\succ{}b\mid{}x)$. We assume that preferences follow the Bradley-Terry model (Bradley and Terry, 1952), in which
@@
for an unknown reward function $r^{\star}:\mathcal{X}\times\mathcal{A}\to\left[0,R_{\mathsf{max}}\right]$ for some $R_{\mathsf{max}}\geq 1$. From the preference dataset $\mathcal{D}_{\mathsf{pref}}$, we aim to learn a policy $\widehat{\pi}$ that has high reward in the sense that
$J(\pi^{\star})-J(\widehat{\pi})\leq\varepsilon,$
for a small $\varepsilon>0$, where $J(\pi)\vcentcolon={}\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x)}\left[r^{\star}(x,a)\right]$ is the true expected reward, and $\pi^{\star}$ is any comparator policy of interest. We abbreviate $\operatorname{\mathbb{E}}_{\pi}\left[\cdot\right]\vcentcolon={}\operatorname{\mathbb{E}}_{x\sim\rho,a\sim\pi(\cdot\mid{}x)}[\cdot]$, and assume that $\rho(x)>0$ for all $x$ and $\pi_{\mathsf{ref}}(a\mid x)>0$ for all $x,a$ without loss of generality.
@@
where we adopt the shorthand $D_{\mathsf{KL}}\left(\pi\,\|\,\pi^{\prime}\right)=\operatorname{\mathbb{E}}_{x\sim\rho}\left[D_{\mathsf{KL}}\left(\pi(\cdot\mid{}x)\,\|\,\pi^{\prime}(\cdot\mid{}x)\right)\right]$. These methods first estimate a reward function $\widehat{r}$ from $\mathcal{D}_{\mathsf{pref}}$ using maximum likelihood under the Bradley-Terry model:
@@
where $\sigma(x)\vcentcolon={}\frac{\exp(x)}{1+\exp(x)}$ is the sigmoid function and $\mathcal{R}$ is a class of reward functions, which is typically parameterized by a neural network. Then, they apply standard policy optimization methods like PPO to optimize an estimated version of the KL-regularized objective:
@@ 2.2 Overoptimization and Insufficiency of KL-Regularization
[…]
@@
In offline reinforcement learning theory, the sample efficiency of an algorithm refers to the number of samples required to guarantee that $J(\widehat{\pi})\approx J(\pi^{\star})$. It is typically quantified by a coverage coefficient (or concentrability coefficient) that measures the quality of the data collected by the reference $\pi_{\mathsf{ref}}$ (Farahmand et al., 2010; Xie and Jiang, 2020; Zanette et al., 2021). We will utilize the $L_{1}$ coverage coefficient, defined for a policy $\pi$ as $\mathcal{C}^{\pi}\vcentcolon={}\operatorname{\mathbb{E}}_{\pi}\left[\frac{\pi(a\mid{}x)}{\pi_{\mathsf{ref}}(a\mid{}x)}\right]$. Single policy concentrability is the gold standard for sample efficiency, and is obtained by an algorithm if, for any comparator policy $\pi^{\star}$, the sample size required to learn $J(\widehat{\pi})\approx J(\pi^{\star})$ scales with $\mathcal{C}^{\pi^{\star}}$, the coverage coefficient of $\pi^{\star}$. This guarantees that $\widehat{\pi}$ is competitive with the best policy that is sufficiently covered by offline data, and, importantly, also guarantees that $\widehat{\pi}$ is never much worse than $\pi_{\mathsf{ref}}$ itself. Single policy concentrability is typically achieved by pessimistic algorithms that penalize the evaluations of candidate policies according to their uncertainty under the offline data, which prevents the learner from overfitting to inaccurate offline reward models.
In contrast, the performance of non-pessimistic algorithms typically scales with all-policy concentrability—meaning that sample complexity scales with $\max_{\pi\in\Pi}\mathcal{C}^{\pi}$ (Liu et al., 2020; Jin et al., 2021; Rashidinejad et al., 2021)— which is a guarantee achieved by even greedy algorithms that directly optimize the offline reward model without regularization. All-policy concentrability describes algorithms that cannot adapt to the quality of the data, and are thereby prone to overoptimization (of the offline reward model) unless the data is rich enough to cover all candidate policies sufficiently well. In contrast, single policy concentrability serves as a theoretical certification that an algorithm is robust to poor data coverage and will not overfit.
@@
Zhu et al. (2023) show that the performance of PPO and DPO scales with all-policy concentrability, $\max_{\pi}\mathcal{C}_{\infty}^{\pi}$, for the stylized case of alignment with linearly parameterized policies where $\pi_{\theta}(a\mid{}x)\propto{}\exp(\left\langle\phi(x,a),\theta\right\rangle)$ for a known feature embedding $\phi(x,a)\in\mathbb{R}^{d}$ (see also Zhu et al. (2024); Song et al. (2024)). They also propose a pessimistic algorithm that achieves
@@
simultaneously for all $\pi^{\star}$. [^1] Zhu et al. (2023) achieve guarantees based on a feature coverage coefficient, which improves on concentrability coefficients, but is specialized to linear function approximation. While encouraging, these results are restricted to linearly parameterized policies, and cannot be directly applied to large language models. Most existing theoretical algorithms for offline alignment are similar in nature, and either place restrictive assumptions on the policy class $\Pi$ (Zhu et al., 2023; Zhan et al., 2023a; Li et al., 2023; Xiong et al., 2023) or are not feasible to implement in a way that is faithful to theory (Ye et al., 2024; Ji et al., 2024).
@@
which augments the DPO objective (the second term) with an additional supervised fine-tuning-like (SFT) loss (the first term). While this objective is simple to apply to general policy classes, the existing single-policy concentrability guarantees for this method assume that $\Pi$ satisfies restrictive convexity conditions which do not hold in practice for large language models. Perhaps surprisingly, we show (Section A.1) that without convexity, the objective in Eq. 5 fails to achieve a single-policy concentrability guarantee. [^2] This finding is rather surprising because Xie et al. (2024) show that an optimistic online counterpart to Eq. 5, which negates the SFT term, enjoys online RLHF guarantees with general policy classes without requiring analogous convexity conditions. In other words, DPO+SFT is insufficient to mitigate overoptimization.